Order: Cefadyl qid IM for a child who weighs 32 lb. The drug book recommends Cefadyl...

  1. Order: Cefadyl qid IM for a child who weighs 32 lb.
    1. The drug book recommends Cefadyl 40 mg/kg/day in divided doses
    2. Give 145 mg per dose.
    3. Cefadyl is provided as 250 mg / 2.5 ml
    4. How many mL would you give?  
Answer #1

Ans) Cefadyl qid IM for a child who weighs 32 lb.
a. The drug book recommends Cefadyl 40 mg/kg/day in divided doses:

lb to kg: 32/2.2 = 14.54 kg.

40 mg/kg/day = 40 × 14.54 = 581.6 ~ 582 mg

Dividing the dose: 582 ÷ 145 = 3.779 ~ 3.8 ml

b. 3.8 ml tds dose i.e 145 mg per dose.

c. Cefadyl is provided as 250 mg / 2.5 ml

250 = 2.5

145 = x

x = 1.45 ~ 1.5 ml.

d. To administer 1.5 ml dose.